1) If an economy can raise its annual real GDP growth rate from 3.8 percent to 4.5 percent, its real GDP doubling time is reduced by 15 years.

2) Suppose that the government passes a law requiring households to increase savings 10% above previous levels. According to Solow's growth theory, in the long run output per capita will grow less rapidly.

3) If K = 3000, n = 0.02, and depreciation, δ= 0.04 and g=0.03, then investment of 320 will hold (K/AL) constant.




1
Expert's answer
2021-08-03T11:49:49-0400

Solution:

1.). The statement is false.

This is because its real GDP doubling time will be reduced by only 3 years.

The rule of 70 is used in estimating the time it takes to double a number based on its growth rate.

Number of years to double real GDP = "\frac{70}{Growth\; rate}"

At 3.8% = "\frac{70}{3.8} = 18 years"


At 4.5% = "\frac{70}{4.5} = 15 years"

Reduction =             3 years

 

2.). The statement is false.

According to Solow’s growth theory, in the long run, output per capita will increase at a higher rate, since a higher saving rate will result in higher steady-state capital stock and a higher level of output.

 

3.). The statement is false.

The investment of 320 will not hold K/AL constant but it will be decreasing.

To breakeven investment:

(n+g+δ)k = s

(0.02+0.03+0.04)3000 = 320

(0.09)3000 = 320

270 < 320